LAB 7 DETERMINATION OF THE MOLAR VOLUME OF A GAS AND THE GAS CONSTANT. R PRE-LAB QUESTIONS 1. A student did the experiment described except she used Al foil instead of Mg. Her data table included the following values. 0.0352 g 19.08 C Mass of Al Temperature Barometric pressure Volume of H, gas Answer the following questions based on this data table. 762.5 mm Hg 46.92 mL a. Write the balanced equation for the reaction between Aluminum and hydrochloric acid. b. How many moles of Al were consumed? c. How many moles of H, were produced? d. What was the vapor pressure of the water in the gas sample?
Answer #1

(a) Balanced equation is

2 Al + 6 HCl --> 2 AlCl3 + 3 H2

2 mol 6 mol 2 mol 3 mol

(c) Given

Volume of H2 gas = 46.92 ml = 0.04692 L

Barometric pressure P = 762.5 mmHg = 762.5 mmHg / 760 = 1.003 atm  

Temperature T = 19.08 C = 292.08 K

R = 0.08206 L.atm/mol.K

PV = nRT

1.003 atm * 0.04692 L = n * 0.08206 L.atm/mol.K * 292.08 K

n = 1.964 * 10-3 mol of H2 Answer (c)

(b) according to reaction stoichiometry

2 moles of Al will be consumed when 3 moles of H2 is produced

No. of moles Al = 1.964 * 10-3 mol * 2 / 3 = 1.31 * 10-3 mol